OpenStudy (anonymous):

what is the vertical asymptote of 3x^2-5x/(3x^2+x-10)

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

= (3x^2 - 5)/(3x-5)(x+2) = x/(x+2) = x = -2 is the vertical asymptote

OpenStudy (mathmate):

Do note that the point x=5/3 is not in the domain of the function.